FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF THE HOSO RADICAL

\underline{VALERIO LATTANZI}, M.C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and School of Engineering and Applied Science, Harvard University, Cambridge, MA 02138; and FILIPPO TAMASSIA, Dipartimento di Chimica Fisica e Inorganica, Università di Bologna, V.le Risorgimento 4, I-40136 Bologna, Italy.

Molecules closely-related to SO_{2} are of considerable interest because of the important role this triatomic or its surrogates play in the atmospheric and combustion chemistry of sulfur-containing compounds. Previous coupled-cluster quantum chemical calculations^{a} of the [HSO_{2}] potential energy surface conclude that a cis-isomer is the ground state, the trans isomer is a transition state, and that a C_{2v} HSO_{2} isomer lies considerably higher in energy (∼ 20 kcal/mol). Using Fourier Transform microwave spectroscopy, the rotational spectrum of the cis-HOSO radical has been precisely characterized for the first time by applying a DC discharge to a mixture of water and sulfur dioxide heavily diluted in neon. The fundamental rotational transition was detected near 17 GHz, in good agreement (within ∼ 1%) of the calculated value derived from the equilibrium structure^{b}. Microwave-microwave double resonance techniques have been used to extend the dataset to higher frequencies and to provide confirmation of the spectroscopic assignments.

aD. Binns, and P. Marshall, Journal of Chemical Physics \underline{95}, 4940, (1991)

bB. Napolion, and J. D. Watts, Chemical Physics Letters \underline{421}, 562, (2006)